### v2

@@ Acknowledgments
We thank Haoxuan Chen, Zhiyan Ding, Michał Dereziński, Jamie Haddock, Jiang Hu, Lin Lin, Anna Ma, Christopher Musco, Deanna Needell, Kevin Shu, Joel Tropp, Roman Vershynin, and Jonathan Weare for helpful conversations.
@@ Appendix A TARK with increasing burn-in time
To control both bias and variance, we recommend implementing TARK with a burn-in time that comprises a quarter to half of the final time, $t_{\rm b}\in[t/4,t/2]$. In practice, we may not know the final time $t$ in advance, opting to run the algorithm for as many iterations as needed for the solution to meet some error tolerance. To implement TARK in a storage-efficient manner in this setting, one can use the following storage-efficient TARK implementation that implements an increasing burn-in time $t_{\rm b}=2^{\lfloor\log_{2}(t)\rfloor-1}$. The approach is based on storing just two extra vectors $\boldsymbol{\widetilde{x}}_{\rm old},\boldsymbol{\widetilde{x}}_{\rm new}\in{}^{d}$, where $\boldsymbol{\widetilde{x}}_{\rm old}$ sums the TARK iterates from time $2^{\lfloor\log_{2}(t)\rfloor-1}$ to $2^{\lfloor\log_{2}(t)\rfloor}$ and $\boldsymbol{\widetilde{x}}_{\rm new}$ sums the TARK iterates from time $2^{\lfloor\log_{2}(t)\rfloor}$ to $t$. When the iteration time $t$ hits a power of $2$, the two vectors are updated via $\boldsymbol{\widetilde{x}}_{\rm old}\leftarrow\boldsymbol{\widetilde{x}}_{\rm new}$ and $\boldsymbol{\widetilde{x}}_{\rm new}\leftarrow\boldsymbol{0}$. At any time $t$, the TARK result can be quickly calculated using $\boldsymbol{\widetilde{x}}_{\rm old}$ and $\boldsymbol{\widetilde{x}}_{\rm new}$ as follows:
$\boldsymbol{\boldsymbol{\widetilde{x}}}_{t}=\frac{1}{t-t_{\rm b}}\sum_{s=t_{\rm b}}^{t-1}\boldsymbol{x}_{s}=\frac{\boldsymbol{\widetilde{x}}_{\rm old}+\boldsymbol{\widetilde{x}}_{\rm new}}{t-t_{\rm b}}.$
@@
2:$\boldsymbol{\widetilde{x}}_{\rm old}=\boldsymbol{0}$, $\boldsymbol{\widetilde{x}}_{\rm new}=\boldsymbol{0}$
@@
5:     $\boldsymbol{x}_{s+1}=\boldsymbol{x}_{s}+\frac{b_{i}-\boldsymbol{a}_{i}^{\top}\boldsymbol{x}_{s}}{\left\|\boldsymbol{a}_{i}\right\|^{2}}\boldsymbol{a}_{i}$
6:     $\boldsymbol{\widetilde{x}}_{\rm new}=\boldsymbol{\widetilde{x}}_{\rm new}+\boldsymbol{x}_{s+1}$
@@
8:         $\boldsymbol{\widetilde{x}}_{\rm old}=\boldsymbol{\widetilde{x}}_{\rm new}$, $\boldsymbol{\widetilde{x}}_{\rm new}=\boldsymbol{0}$
@@
12:$\boldsymbol{\overline{x}}_{t}=\frac{1}{t-t_{\rm b}}(\boldsymbol{\widetilde{x}}_{\rm old}+\boldsymbol{\widetilde{x}}_{\rm new})$
13:return $\boldsymbol{\overline{x}}_{t}$
@@ Appendix B Lower bounds
Proposition B.1 (Lower bound on mean square error).
Fix $\varepsilon>0$ and $d\geq 1$. Any algorithm that can solve all least-squares problems involving all matrices $\boldsymbol{A}\in\mathbb{R}^{n\times d}$ and vectors $\boldsymbol{b}\in\mathbb{R}^{n}$ with mean square error
(9) $\operatorname{\mathbb{E}}\lVert\widehat{\boldsymbol{x}}-\boldsymbol{x}_{\star}\rVert^{2}\leq\varepsilon\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$
@@
Consider applying any least-squares solver to a random class of least-squares problems $\min_{\boldsymbol{x}}\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\right\|^{2}$ of the form
@@
$\boldsymbol{A}=\begin{bmatrix}\boldsymbol{1}&\cdots&\boldsymbol{0}\\ \vdots&\ddots&\vdots\\ \boldsymbol{0}&\cdots&\boldsymbol{1}\end{bmatrix}\in{}^{md\times d}\quad\text{and}\quad\boldsymbol{b}=\begin{bmatrix}\boldsymbol{b}_{1}\\ \vdots\\ \boldsymbol{b}_{d}\end{bmatrix},$
where $\boldsymbol{0},\boldsymbol{1}\in{}^{m}$ are the vectors of all zeroes and all ones, respectively. Each problem decomposes into the sum of $d$ simpler problems $\min_{\boldsymbol{x}\in{}^{d}}\sum\nolimits_{i=1}^{d}\lVert\boldsymbol{b}_{i}-\boldsymbol{1}x_{i}\rVert^{2}$. The random noise comes from the vectors $\boldsymbol{b}_{1},\ldots,\boldsymbol{b}_{d}$, which are generated according to
$\boldsymbol{b}_{i}=\boldsymbol{1}y_{i}+\boldsymbol{z}_{i},\quad\text{for }\begin{cases}y_{i}\sim\mathcal{N}(0,\sigma^{2}),\\ \boldsymbol{z_{i}}\sim\mathcal{N}\bigl{(}\boldsymbol{0},\mathbf{I}_{m}-\tfrac{1}{m}\boldsymbol{1}\boldsymbol{1}^{\top}\bigr{)}.\end{cases}$
In this setup, observe that the joint distribution of $\boldsymbol{b}_{i}$ and $y_{i}$ is given by
(10) $\begin{bmatrix}\boldsymbol{b_{i}}\\ y_{i}\end{bmatrix}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma}),\quad\text{where }\boldsymbol{\Sigma}=\begin{bmatrix}\sigma^{2}+1-\tfrac{1}{m}&\cdots&\sigma^{2}-\tfrac{1}{m}&\sigma^{2}\\ \vdots&\ddots&\vdots&\vdots\\ \sigma^{2}-\tfrac{1}{m}&\cdots&\sigma^{2}+1-\tfrac{1}{m}&\sigma^{2}\\ \sigma^{2}&\cdots&\sigma^{2}&\sigma^{2}\end{bmatrix}.$
Also, the mean and covariance formulas for $\boldsymbol{z}_{i}$ show that $\boldsymbol{1}^{\top}\boldsymbol{z}_{i}=0$ almost surely. Hence, the optimal least-squares error comes from setting $\boldsymbol{x}_{\star}=\begin{bmatrix}y_{1}&\cdots&y_{d}\end{bmatrix}^{\top}$, which leads to
@@
(11) $\operatorname{\mathbb{E}}\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\right\|^{2}=\sum\nolimits_{i=1}^{d}\operatorname{\mathbb{E}}\left\|\boldsymbol{z}_{i}\right\|^{2}=d(m-1).$
This is the expected square error of the optimal least-squares solution.
@@
Now suppose a least-squares solver only accesses a subset of the entries of $\boldsymbol{b}_{1}$ with index set $\mathsf{S}_{1}\subseteq\{1,\ldots,m\}$, a subset of the entries of $\boldsymbol{b}_{2}$ with index set $\mathsf{S}_{2}\subseteq\{1,\ldots,m\}$, and so on. Conditional on the revealed entries, the vector $\bigl{[}\begin{smallmatrix}\boldsymbol{b}_{i}\\ y_{i}\end{smallmatrix}\bigr{]}$ now has a Gaussian distribution. In particular, applying the mean and variance formulas for Gaussian conditioning yields $y_{i}\sim\mathcal{N}(m_{i},v_{i}^{2})$ for
@@
$m_{i}=\boldsymbol{\Sigma}(m+1,\mathsf{S}_{i})\boldsymbol{\Sigma}(\mathsf{S}_{i},\mathsf{S}_{i})^{-1}\boldsymbol{b}_{i}(\mathsf{S}_{i})=\frac{\sigma^{2}}{|\textsf{S}_{i}|\bigl{(}\sigma^{2}-\frac{1}{m}\bigr{)}+1}\sum_{j\in\mathsf{S}_{i}}b_{j},$
@@
$v_{i}^{2}=[\boldsymbol{\Sigma}]_{m+1,m+1}-\boldsymbol{\Sigma}(m+1,\mathsf{S}_{i})\boldsymbol{\Sigma}(\mathsf{S}_{i},\mathsf{S}_{i})^{-1}\boldsymbol{\Sigma}(\mathsf{S}_{i},m+1)=\sigma^{2}-\frac{\sigma^{4}|\mathsf{S}_{i}|}{|\textsf{S}_{i}|\bigl{(}\sigma^{2}-\frac{1}{m}\bigr{)}+1}.$
These equations can be verified by consulting the formula for $\boldsymbol{\Sigma}$ in Eq. 10. Conditional on the revealed entries of $\boldsymbol{b}$, the conditional mean $m_{i}$ is the optimal estimator of $y_{i}$:
@@
$m_{i}=\operatornamewithlimits{argmin}_{\widehat{x}_{i}}\operatorname{\mathbb{E}}\bigl{[}|\widehat{x}_{i}-y_{i}|^{2}\,|\,\boldsymbol{b}_{i}(\mathsf{S}_{i})\bigr{]}.$
@@
Averaging over the randomness in the unrevealed entries of $\boldsymbol{b}$, it holds for any estimator $\boldsymbol{\widehat{x}}$
@@
$\operatorname{\mathbb{E}}\bigl{[}\lVert\boldsymbol{\widehat{x}}-\boldsymbol{x}_{\star}\rVert^{2}\,|\,\boldsymbol{b}_{1}(\mathsf{S}_{1}),\ldots,\boldsymbol{b}_{d}(\mathsf{S}_{d})\bigr{]}$ $\geq\operatorname{\mathbb{E}}\bigl{[}\lVert\boldsymbol{m}-\boldsymbol{x}_{\star}\rVert^{2}\,|\,\boldsymbol{b}_{1}(\mathsf{S}_{1}),\ldots,\boldsymbol{b}_{d}(\mathsf{S}_{d})\bigr{]}$
$=\sum_{i=1}^{d}v_{i}^{2}=d\sigma^{2}-\sum_{i=1}^{d}\frac{\sigma^{4}|\mathsf{S}_{i}|}{|\mathsf{S}_{i}|(\sigma^{2}-\frac{1}{m})+1}.$
Further, averaging over the randomness in all the entries of $\boldsymbol{b}$,
$\operatorname{\mathbb{E}}\lVert\boldsymbol{\widehat{x}}-\boldsymbol{x}_{\star}\rVert^{2}\geq\operatorname{\mathbb{E}}\Biggl{[}d\sigma^{2}-\sum_{i=1}^{d}\frac{\sigma^{4}|\mathsf{S}_{i}|}{|\mathsf{S}_{i}|(\sigma^{2}-\frac{1}{m})+1}\Biggr{]}.$
@@
To simplify this mean square error formula, observe that as long as $\sigma^{2}>1/m$, the function $|\mathsf{S}_{i}|\mapsto-\sigma^{4}|\mathsf{S}_{i}|/\bigl{[}|\mathsf{S}_{i}|(\sigma^{2}-\frac{1}{m})+1\bigr{]}$ is decreasing and convex. Since $\frac{1}{d}\sum_{i=1}^{d}|\mathsf{S}_{i}|\leq t/d$, where $t$ is the maximum number of entries accessed, it follows
@@
$\operatorname{\mathbb{E}}\lVert\boldsymbol{\widehat{x}}-\boldsymbol{x}_{\star}\rVert^{2}\geq d\sigma^{2}-\sum_{i=1}^{d}\frac{\sigma^{4}\cdot\frac{t}{d}}{\frac{t}{d}(\sigma^{2}-\frac{1}{m})+1}=d\sigma^{2}-\frac{d\sigma^{4}}{\sigma^{2}-\frac{1}{m}+\frac{d}{t}}=\biggl{[}\frac{d^{2}}{t}-\frac{d}{m}\biggr{]}\cdot\frac{\sigma^{2}}{\sigma^{2}-\frac{1}{m}+\frac{d}{t}}.$
This lower bound constrains the accuracy of the least-squares solver when applied to the random problem class.
Now, suppose that the least-squares solver satisfies Eq. 9. Then for this problem class the approximation error must satisfy
$\operatorname{\mathbb{E}}\lVert\widehat{\boldsymbol{x}}-\boldsymbol{x}_{\star}\rVert^{2}\leq\varepsilon\cdot\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\operatorname{\mathbb{E}}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}=\varepsilon d\cdot\frac{m-1}{m},$
since $\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}=1/m$ and $\operatorname{\mathbb{E}}\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}=m(d-1)$ Eq. 11. Last, the relation
$\biggl{[}\frac{d^{2}}{t}-\frac{d}{m}\biggr{]}\cdot\frac{\sigma^{2}}{\sigma^{2}-\frac{1}{m}+\frac{d}{t}}\leq\varepsilon d\cdot\frac{m-1}{m}$
can only hold for arbitrarily large values of $\sigma^{2}$ and $m$ if the maximum number of entries accessed is $t\geq d/\varepsilon$.
Proposition B.1 also leads to a bound on the mean square residual error $\operatorname{\mathbb{E}}\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}$.
Corollary B.3 (Lower bound on mean square residual error).
Fix $\varepsilon>0$ and $d\geq 1$. Any algorithm that can solve all least-squares problems involving all matrices $\boldsymbol{A}\in\mathbb{R}^{m\times d}$ and vectors $\boldsymbol{b}\in\mathbb{R}^{m}$ with mean square residual error
@@
(12) $\operatorname{\mathbb{E}}\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}\leq(1+\varepsilon)\cdot\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$

### v3

@@ Acknowledgments
The authors thank Haoxuan Chen, Zhiyan Ding, Michał Dereziński, Jamie Haddock, Jiang Hu, Lin Lin, Anna Ma, Christopher Musco, Deanna Needell, Kevin Shu, Joel Tropp, Roman Vershynin, and Jonathan Weare for helpful conversations. ENE acknowledges support from the Department of Energy under Award Number DE-SC0021110 and, under the aegis of Joel Tropp, from NSF FRG 1952777 and the Carver Mead New Horizons Fund. GG acknowledges support from the Department of Energy under Award Number DE-SC0023112.
@@ Appendix A TARK with increasing burn-in time
To control both bias and variance, we recommend implementing TARK with a burn-in time that comprises a quarter to half of the final time, $t_{\rm b}\in[t/4,t/2]$. In practice, we may not know the final time $t$ in advance, opting to run the algorithm for as many iterations as needed for the solution to meet some error tolerance. To implement TARK in a storage-efficient manner in this setting, one can use the following TARK implementation with an increasing burn-in time $t_{\rm b}=2^{\lfloor\log_{2}(t)\rfloor-1}$. The approach is based on storing just two extra vectors $\widetilde{\boldsymbol{x}}_{\rm old},\widetilde{\boldsymbol{x}}_{\rm new}\in{}^{d}$, where $\widetilde{\boldsymbol{x}}_{\rm old}$ sums the TARK iterates from time $2^{\lfloor\log_{2}(t)\rfloor-1}$ to $2^{\lfloor\log_{2}(t)\rfloor}$ and $\widetilde{\boldsymbol{x}}_{\rm new}$ sums the TARK iterates from time $2^{\lfloor\log_{2}(t)\rfloor}$ to $t$. When the iteration time $t$ hits a power of $2$, the two vectors are updated via $\widetilde{\boldsymbol{x}}_{\rm old}\leftarrow\widetilde{\boldsymbol{x}}_{\rm new}$ and $\widetilde{\boldsymbol{x}}_{\rm new}\leftarrow\boldsymbol{0}$. At any time $t$, the TARK result can be quickly calculated using $\widetilde{\boldsymbol{x}}_{\rm old}$ and $\widetilde{\boldsymbol{x}}_{\rm new}$ as follows:
$\widetilde{\boldsymbol{\boldsymbol{x}}}_{t}=\frac{1}{t-t_{\rm b}}\sum\nolimits_{s=t_{\rm b}}^{t-1}\boldsymbol{x}_{s}=\frac{\widetilde{\boldsymbol{x}}_{\rm old}+\widetilde{\boldsymbol{x}}_{\rm new}}{t-t_{\rm b}}.$
@@
2:$\widetilde{\boldsymbol{x}}_{\rm old}=\boldsymbol{0}$, $\widetilde{\boldsymbol{x}}_{\rm new}=\boldsymbol{0}$
@@
5:     $\boldsymbol{x}_{s+1}=\boldsymbol{x}_{s}+(b_{i}-\boldsymbol{a}_{i}^{\top}\boldsymbol{x}_{s})\,\boldsymbol{a}_{i}/\left\|\boldsymbol{a}_{i}\right\|^{2}$
6:     $\widetilde{\boldsymbol{x}}_{\rm new}=\widetilde{\boldsymbol{x}}_{\rm new}+\boldsymbol{x}_{s+1}$
@@
8:         $\widetilde{\boldsymbol{x}}_{\rm old}=\widetilde{\boldsymbol{x}}_{\rm new}$, $\widetilde{\boldsymbol{x}}_{\rm new}=\boldsymbol{0}$
@@
12:$\overline{\boldsymbol{x}}_{t}=(\widetilde{\boldsymbol{x}}_{\rm old}+\widetilde{\boldsymbol{x}}_{\rm new})/(t-t_{\rm b})$
13:return $\overline{\boldsymbol{x}}_{t}$
@@ Appendix B Lower bounds
Proposition 1 (Lower bound on mean square error).
Fix $\varepsilon>0$ and $d\geq 1$. Any algorithm that can solve all least-squares problem $\min_{\boldsymbol{x}}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\rVert^{2}$ with mean square error
$\operatorname{\mathbb{E}}\,\lVert\widehat{\boldsymbol{x}}-\boldsymbol{x}_{\star}\rVert^{2}\leq\varepsilon\,\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2},$ (9)
@@
Consider applying a least-squares solver to a a random least-squares problem, where
@@
$\boldsymbol{A}=\begin{bmatrix}\boldsymbol{1}_{m}&\cdots&\boldsymbol{0}_{m}\\ \vdots&\ddots&\vdots\\ \boldsymbol{0}_{m}&\cdots&\boldsymbol{1}_{m}\end{bmatrix}\in{}^{md\times d},\quad\boldsymbol{b}=\begin{bmatrix}\boldsymbol{b}_{1}\\ \vdots\\ \boldsymbol{b}_{d}\end{bmatrix}\in{}^{md},\quad\text{and }\boldsymbol{b}_{i}\sim\text{iid }\mathcal{N}\bigl{(}\boldsymbol{0},\boldsymbol{\Sigma}\bigr{)},\text{ for }\boldsymbol{\Sigma}=\mathbf{I}_{m}+v\boldsymbol{1}_{m}\boldsymbol{1}_{m}^{\top}.$ (10)
Here, $\boldsymbol{0}_{m},\boldsymbol{1}_{m}\in{}^{m}$ are the vectors of all zeroes and all ones, $m$ controls the aspect ratio in the problem, and $v>0$ is a variance parameter. Each problem decomposes into the sum of $d$ simpler problems:
$\min_{\boldsymbol{x}\in{}^{d}}\,\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\right\|^{2}=\min_{\boldsymbol{x}\in{}^{d}}\sum\nolimits_{i=1}^{d}\lVert\,\boldsymbol{b}_{i}-\boldsymbol{1}_{m}x_{i}\,\rVert^{2}.$
Hence, each entry of the least-squares solution $\boldsymbol{x}_{\star}\in{}^{d}$ is an arithmetic mean of $m$ entries of the output vector:
$\boldsymbol{x}_{\star}=\begin{bmatrix}x_{\star,1}&\cdots&x_{\star,d}\end{bmatrix}^{\top}=\begin{bmatrix}\frac{1}{m}\boldsymbol{1}_{m}^{\top}\boldsymbol{b}_{1}&\cdots&\frac{1}{m}\boldsymbol{1}_{m}^{\top}\boldsymbol{b}_{d}\end{bmatrix}^{\top}.$
In this random problem class, a direct calculation using the Gaussian covariance matrix $\boldsymbol{\Sigma}=\mathbf{I}_{m}+v\boldsymbol{1}_{m}\boldsymbol{1}_{m}^{\top}$ shows
@@
$\operatorname{\mathbb{E}}\,\left\|\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\right\|^{2}=\sum\nolimits_{i=1}^{d}\operatorname{\mathbb{E}}\,\Bigl{\lVert}\boldsymbol{b}_{i}-\tfrac{1}{m}\boldsymbol{1}_{m}\boldsymbol{1}_{m}^{\top}\boldsymbol{b}_{i}\Bigr{\rVert}^{2}=d(m-1).$
This is the mean square error of the least-squares solution. Further, the matrix $\boldsymbol{A}$ has singular values $\sigma_{i}=\sqrt{m}$ for $i=1,\ldots,m$, so $\lVert\boldsymbol{A}^{+}\rVert^{2}=\frac{1}{m}$.
@@
Now suppose a least-squares solver accesses certain entries of $\boldsymbol{b}_{1}$ that are indexed by $\mathsf{S}_{1}\subseteq\{1,\ldots,m\}$, certain entries of $\boldsymbol{b}_{2}$ that are indexed by $\mathsf{S}_{2}\subseteq\{1,\ldots,m\}$, and so on. Given a subset of $k=|\textsf{S}_{i}|$ revealed entries of $\boldsymbol{b}_{i}$, evaluate the Gaussian conditional mean and variance formulas for the unrevealed entries $\boldsymbol{b}_{i,\textsf{S}_{i}^{c}}$ as follows:
$\mathbb{E}\,\bigl{[}\boldsymbol{b}_{i,\textsf{S}_{i}^{c}}\,\big{|}\,\boldsymbol{b}_{i,\textsf{S}_{i}}\bigr{]}=\boldsymbol{\Sigma}_{\textsf{S}_{i}^{c},\textsf{S}_{i}}\,\boldsymbol{\Sigma}_{\textsf{S}_{i},\textsf{S}_{i}}^{-1}\,\boldsymbol{b}_{i,\textsf{S}_{i}}=\frac{v}{1+vk}\boldsymbol{1}_{m-k}\boldsymbol{1}_{k}^{\top}\boldsymbol{b}_{i,\textsf{S}_{i}}.$
and also
$\operatorname{cov}\,\bigl{[}\boldsymbol{b}_{i,\textsf{S}_{i}^{c}}\,|\,\boldsymbol{b}_{i,\textsf{S}_{i}}]=\boldsymbol{\Sigma}_{\textsf{S}_{i},\textsf{S}_{i}}-\boldsymbol{\Sigma}_{\textsf{S}_{i}^{c},\textsf{S}_{i}}\,\boldsymbol{\Sigma}_{\textsf{S}_{i},\textsf{S}_{i}}^{-1}\,\boldsymbol{\Sigma}_{\textsf{S}_{i},\textsf{S}_{i}^{c}}=\mathbf{I}_{m-k}+\frac{v}{1+vk}\boldsymbol{1}_{m-k}\boldsymbol{1}_{m-k}^{\top}.$
Therefore, the conditional expectation formula for $x_{\star,i}$ is
@@
$\mathbb{E}\,\bigl{[}x_{\star,i}\,\big{|}\,\boldsymbol{b}_{i,\textsf{S}_{i}}\bigr{]}=\mathbb{E}\,\biggl{[}\frac{1}{m}\boldsymbol{1}_{m}^{\top}\boldsymbol{b}_{i}\,\bigg{|}\,\boldsymbol{b}_{i,\textsf{S}_{i}}\biggr{]}=\frac{1}{m}\biggl{[}1+\frac{v(m-k)}{1+vk}\biggr{]}\boldsymbol{1}_{k}^{\top}\boldsymbol{b}_{i,\textsf{S}_{i}}.$
The conditional variance formula for $x_{\star,i}$ is
@@
$\operatorname{Var}\,\bigl{[}x_{\star,i}\,\big{|}\,\boldsymbol{b}_{i,\textsf{S}_{i}}\bigr{]}=\operatorname{Var}\,\biggl{[}\frac{1}{m}\boldsymbol{1}_{m}^{\top}\boldsymbol{b}_{i}\,\bigg{|}\,\boldsymbol{b}_{i,\textsf{S}_{i}}\biggr{]}=\frac{\boldsymbol{1}_{m-k}^{\top}\operatorname{cov}\,\bigl{[}\boldsymbol{b}_{i,\textsf{S}_{i}^{c}}\,\big{|}\,\boldsymbol{b}_{i,\textsf{S}_{i}}\bigr{]}\,\boldsymbol{1}_{m-k}}{m^{2}}=\frac{m-k}{m^{2}}\biggl{[}1+\frac{v(m-k)}{1+vk}\biggr{]}$
The vector with entries $\mu_{i}=\mathbb{E}\,\bigl{[}x_{\star,i}\,\big{|}\,\boldsymbol{b}_{i,\textsf{S}_{i}}\bigr{]}$ optimizes the mean square error of approximating $\boldsymbol{x}_{\star}$.
@@
$\mu_{i}=\operatornamewithlimits{argmin}_{\widehat{x}_{i}}\operatorname{\mathbb{E}}\,\bigl{[}|\,\widehat{x}_{i}-x_{\star,i}|^{2}\,|\,\boldsymbol{b}_{i,\mathsf{S}_{i}}\bigr{]}.$
@@
Therefore, taking the expectation over the unrevealed entries of $\boldsymbol{b}$, it holds for any estimator $\widehat{\boldsymbol{x}}$:
@@
$\operatorname{\mathbb{E}}\,\bigl{[}\lVert\,\widehat{\boldsymbol{x}}-\boldsymbol{x}_{\star}\rVert^{2}\,|\,\boldsymbol{b}_{1,\mathsf{S}_{1}},\ldots,\boldsymbol{b}_{d,\mathsf{S}_{d}}\bigr{]}$ $\geq\operatorname{\mathbb{E}}\,\bigl{[}\lVert\,\boldsymbol{\mu}-\boldsymbol{x}_{\star}\rVert^{2}\,|\,\boldsymbol{b}_{1,\mathsf{S}_{1}},\ldots,\boldsymbol{b}_{d,\mathsf{S}_{d}}\bigr{]}$
$=\frac{1}{m^{2}}\sum\nolimits_{i=1}^{d}(m-|\textsf{S}_{i}|)\,\biggl{[}1+\frac{v(m-|\textsf{S}_{i}|)}{1+v|\textsf{S}_{i}|}\biggr{]}.$
@@
Now let $t$ be the maximum number of entries accessed, and observe that $x\mapsto(m-x)\bigl{[}1+v(m-x)/(1+vx)\bigr{]}$ is convex and decreasing, so the conditional mean square error is bounded from below by setting $|\textsf{S}_{i}|=t/d$ for each $i\in\{1,\ldots,d\}$:
@@
$\operatorname{\mathbb{E}}\,\bigl{[}\lVert\widehat{\boldsymbol{x}}-\boldsymbol{x}_{\star}\rVert^{2}\,|\,\boldsymbol{b}_{1,\mathsf{S}_{1}},\ldots,\boldsymbol{b}_{d,\mathsf{S}_{d}}\bigr{]}\geq\frac{d}{m^{2}}(m-t/d)\,\biggl{[}1+\frac{v(m-t/d)}{1+vt/d}\biggr{]}$
By averaging over the revealed entries, the mean square error satisfies the same error bound:
@@
$\operatorname{\mathbb{E}}\,\lVert\widehat{\boldsymbol{x}}-\boldsymbol{x}_{\star}\rVert^{2}\geq\frac{d}{m^{2}}(m-t/d)\,\biggl{[}1+\frac{v(m-t/d)}{1+vt/d}\biggr{]}.$
This is the minimal least-squares solver that an algorithm can possibly achieve after revealing $t$ entries of $\boldsymbol{b}$.
Now suppose that a least-squares solver satisfies the error bound (9) for each linear least-squares problem. Then apply the least-squares solver to the random problem class (10) and average the resulting error bounds (9) to yield
$\frac{d}{m^{2}}(m-t/d)\,\biggl{[}1+\frac{v(m-t/d)}{1+vt/d}\biggr{]}\leq\operatorname{\mathbb{E}}\,\lVert\widehat{\boldsymbol{x}}-\boldsymbol{x}_{\star}\rVert^{2}\leq\varepsilon\,\left\|\smash{\boldsymbol{A}^{+}}\right\|^{2}\,\operatorname{\mathbb{E}}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}=\varepsilon\frac{d(m-1)}{m}.$ (11)
The relation (11) must hold for any parameters $m$ and $v$. Sending $v\rightarrow\infty$ implies that
$t\geq\frac{d^{2}}{\varepsilon d+(1-\varepsilon)\frac{d}{m}}.$
Next, sending $m\rightarrow\infty$ implies that the maximal number of entries which need to be accessed by the algorithm is $t\geq d/\varepsilon$. ∎
Proposition 1 also leads to a bound on the mean square residual error $\operatorname{\mathbb{E}}\,\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}$.
Corollary 1 (Lower bound on mean square residual error).
Fix $\varepsilon>0$ and $d\geq 1$. Any algorithm that can solve all least-squares problem $\min_{\boldsymbol{x}}\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x}\rVert^{2}$ with mean square residual error
$\operatorname{\mathbb{E}}\,\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}\leq(1+\varepsilon)\,\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$ (12)